Bookmatch 8-14 (book/static) Question Help (Measuring growth) Given that a firms return on equity is 18 percent and manageme(Common stock valuation) Sanford common stock is expected to pay $1.85 in dividends next year, and the market price is projec(Common stock valuation) Sanford common stock is expected to pay $1.85 in dividends next year, and the market price is projec(Common stock valuation) The common stock of NCP paid $1.50 in dividends last year. Dividends are expected to grow at an annu(Common stock valuation) Abercrombie & Fitchs common stock pays a dividend of $1.50. It is currently selling for $37.88. If(Common stock valuation) Schlumberger is selling for $64.91 per share and paid a dividend of $1.10 last year. The dividend is
Answer #1

Growth Rate = Return on Equity*Retention Ratio

= 18%*0.4

= 7.2%

Current value = (Expected Dividend+Market price)/(1+Required Return)

= (1.85+51.35)/(1.09)

= $48.81

Value = (1.85+42.50)/(1+0.11)

= $39.95

a.Price = Expected Dividend/(Required Return – Growth rate)

27.81 = 1.50(1+0.0840)/(Expected Return – 8.4%)

Expected Return = 14.25%

Value = 1.50(1+8.4%)/(10.4%-8.4%)

= $81.3

c.Yes, since rate is higher than expected

37.88 = 1.50(1+g)/(16%-g)

6.0608 – 37.88g = 1.50 + 1.50g

G = 11.58%

64.91 = 1.10(1+4%)/(Expected Return – 4%)

Return = 5.76%